OpenStudy (anonymous):

Solve for x: |x + 2| + 16 = 14

OpenStudy (mathmale):

Strong hints: 1) subtract 16 from both sides. 2) look very carefully at what's left of this equation after simplification. Either it can or cannot be solved. Which is the case, and why?

OpenStudy (anonymous):

This problem is very easy. The absolute value symbol doesn't do anything since both terms are constant. The problems then simplifies to x+18=14 -18 -18 x= -4

OpenStudy (anonymous):

Both terms are positive*

OpenStudy (mathmale):

@stefanwolf1: You've come up with x=-4 as a possible solution. Would you please substitute that into the original equation as a check?